OpenStudy (anonymous):

Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain.

OpenStudy (anonymous):

n a_n 1 9 2 3 3 -3

OpenStudy (anonymous):

A. a_n = 9 − 3(n − 1) where n ≤ 9 B. a_n = 9 − 3(n − 1) where n ≥ 1 C. a_n = 9 − 6(n − 1) where n ≤ 9 D. a_n = 9 − 6(n − 1) where n ≥ 1

hartnn (hartnn):

so, you know the formula is an = a1 +(n-1)d a1 = 1st term = 9 can you find the common difference d = ... ?

hartnn (hartnn):

or know what common difference means ?

OpenStudy (anonymous):

Can you walk me through it? I think the common difference is 3..??

hartnn (hartnn):

ok, sure common difference means the difference from next term to current term! so, d = 2nd term - 1st term = 3rd tem - 2nd term = 4th term -3rd term got this ?

OpenStudy (anonymous):

Kinda...

hartnn (hartnn):

here, d = 3-9 = -3 - (3) = ...?

hartnn (hartnn):

let me give you something to read, maybe you'll understand more.... Tutorial on AP,GP,HP. http://openstudy.com/study#/updates/503bb2a0e4b007f900310

hartnn (hartnn):

so what you got d here ?

OpenStudy (anonymous):

-6?

hartnn (hartnn):

d = -6 is correct :)

hartnn (hartnn):

an= a1 +(n-1)d = 9 + (n-1)(-6)

hartnn (hartnn):

so which choice ? also not that 'n' is natural number, because we count it, 1st term, 2nd term and so on...so it has to be n >=1

OpenStudy (anonymous):

a_n=9-6(n-1) where n ≥ 1

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

You're the best! :D

hartnn (hartnn):

thank you for the compliment :) always happy to help :)